If you can give me explanations that would be helpful, thanks.

The following is taken from the Sandhill Company balance sheet. Sandhill Company Balance Sheet (partial) December 31, 2020 CuJournalize the payment of bond interest on January 1, 2021. (Credit account titles are automatically indented when amount isPrepare the entry to amortize bond premium and to accrue the interest due on December 31, 2021. (Credit account titles are auAssume that on January 1, 2022, after paying interest, Sandhill Company calls bonds having a face value of $1,208,000. The caPrepare the adjusting entry at December 31, 2022, to amortize bond premium and to accrue interest on the remaining bonds. (Cr
Answer #1

Solution

Sandhill Company

  1. Entry to record payment of bond interest:

Date

Account Titles and Explanation

Debit

Credit

Jan 1, 2021

Interest Payable

$271,800

Cash

$271,800

(To record payment of interest on Jan 1, 2021)

The interest on bonds is paid on January 1 annually. So, the interest payable at December 31, 2020 is paid on January 1, 2021. The above entry records the interest payment.

  1. Entry to amortize bond premium and accrue the interest due:

Date

Account Titles and Explanation

Debit

Credit

Dec 31, 2021

Interest Expense

$250,600

Premium on bonds payable

$21,200

Interest Payable

$271,800

(To record premium on bonds payable and the accrued interest)

Computations:

Amortization of premium –

Premium on bonds payable = $212,000

Using straight line amortization,

Bonds will be outstanding for 10 years

Annual premium amortization = 212,000/10 = $21,200

Interest expense = interest payable – premium amortization

Interest payable = 9% x 3,020,000 = $271,800

Interest expense = 271,800 – 21,200 = $250,600

  1. Entry to record redemption of the bond:

Date

Account Titles and Explanation

Debit

Credit

Jan 1, 2022

Bonds Payable

$1,208,000

Premium on bonds payable

$76,320

Gain on bond redemption

$28,000

Cash

$1,256,320

(To record redemption of bonds called at $104)

Computations:

Premium on bonds redeemed –

Premium on bonds payable balance, Jan 1, 2022 = 212,000 – 21,200 = $190,800

Premium on bonds redeemed = 1,208,000 x (190,800/3,020,000) = $76,320

Cash received on bonds redeemed = 1,208,000 x 104% = $1,256,320

Gain on redemption = (1,208,000 + 76,320) – 1,256,320 = $28,000

  1. Adjusting entry at December 31, 2022 to amortize bond premium and to accrue bond interest:

Date

Account Titles and Explanation

Debit

Credit

Dec 31, 2022

Interest Expense

$150,360

Premium on bonds payable

$12,720

Interest Payable

$163,080

(To record interest payable and premium amortization)

Interest payable = 9% x (3,020,000 – 1,208,000) = $163,080

Premium amortization = (212,000 – 21,200 – 76,320) x 1/9 years = 114,480/9 = $12,720

Interest expense = 163,080 – 12,720 = $150,360

Note: interest at 9% is payable on the balance in the bonds payable account.

Balance in bonds payable = 3,020,000 – 1,208,000 = 1,812,000

Premium on bonds payable balance = 212,000 – first year amortization, 21,200 – amortization on bonds redeemed, 76,320 = $114,480

Remaining years to bond maturity = 10 -1 = 9 years

Hence, premium of $114,480 is amortized over 9 years.